# T-cell receptor specificity landscape revealed through de novo peptide design

### v1

@@ Abstract
T-cells play a key role in adaptive immunity by mounting specific responses against diverse pathogens. An effective binding between T-cell receptors (TCRs) and pathogen-derived peptides presented on Major Histocompatibility Complexes (MHCs) mediate an immune response. However, predicting these interactions remains challenging due to limited functional data on T-cell reactivities. Here, we introduce a computational approach to predict TCR interactions with peptides presented on MHC class I alleles, and to design novel immunogenic peptides for specified TCR-MHC complexes. Our method leverages HERMES, a structure-based, physics-guided machine learning model trained on the protein universe to predict amino acid preferences based on local structural environments. Despite no direct training on TCR-pMHC data, the implicit physical reasoning in HERMES enables us to make accurate predictions of both TCR-pMHC binding affinities and T-cell activities across diverse viral epitopes and cancer neoantigens, achieving up to 72% correlation with experimental data. Leveraging our TCR recognition model, we develop a computational protocol for de novo design of immunogenic peptides. Through experimental validation in three TCR-MHC systems targeting viral and cancer peptides, we demonstrate that our designs—with up to five substitutions from the native sequence—activate T-cells at success rates of up to 50%. Lastly, we use our generative framework to quantify the diversity of the peptide recognition landscape for various TCR-MHC complexes, offering key insights into T-cell specificity in both humans and mice. Our approach provides a platform for immunogenic peptide and neoantigen design, opening new computational paths for T-cell vaccine development against viruses and cancer.
@@ I Introduction
Predicting the TCR-pMHC binding specificity is an important step in characterizing the efficacy of the immune system in countering different pathogens. Such understanding would aid with disease diagnoses and development of diagnostic tests for early detection of autoimmune diseases [1]. A predictive model for TCR-pMHC specificity can be used to design engineered TCRs that recognize cancer antigens [2], enhancing the effectiveness of adoptive cell transfer and CAR T-cell therapies [3, 4], as well as the design of soluble TCRs and bispecific T-cell engager therapeutics [5]. Moreover, it can enable the design of optimized antigens for vaccines to elicit robust T-cell responses against specific pathogens or tumors [6, 7].
@@
One bottleneck in deciphering the TCR-pMHC code is the lack of well-curated datasets. High-throughput immune repertoire sequencing, combined with comparative analyses of hosts facing similar immune challenges, and the development of functional assays (e.g., pMHC-tetramer staining [8] or MIRA [9]), have enabled the identification of diverse TCR pools with potential reactivity to specific antigens [10], collected in repositories like Immune Epitope Database (IEDB) [11] and VDJdb [12]. However, the current databases are unbalanced, with only a handful of dominant peptides (e.g. derived from SARS-CoV-2, CMV, EBV, Influenza, etc.) and MHC alleles present, each linked to hundreds of TCRs. As a result, statistical inference and machine learning methods trained on such data to predict TCR-pMHC reactivities lack generalizability beyond their training data [13, 14, 15]. On the structural side, only a few hundred TCR-pMHC crystal structures have been experimentally generated to date. This is in part because TCR-pMHC interactions are typically of low affinity and are difficult to stabilize to the level needed for CryoEM or X-ray crystallography [16]. Importantly, this scarcity hinders computational prediction methods like AlphaFold in modeling reliable TCR-pMHC structures [17, 18].
@@
Machine learning models, such as large language models, pre-trained on vast unannotated protein datasets can generate meaningful data representations [19, 20]. Fine-tuning these models on smaller, labeled datasets for specific tasks (e.g., predicting mutation impacts on stability or function) has proven highly effective [21, 22, 23, 24]. Pre-trained sequence-based protein language models have been employed to predict reactivities of antibodies to antigens [23, 24], or TCRs to pMHC complexes [25, 15]. However, lack of generalizability due to sparse and skewed training data has remained a challenge in these models.
@@
Figure 1: HERMES model for predicting and designing TCR-pMHC interactions. (A) HERMES is a structure-based, rotationally equivariant neural network pre-trained on the protein universe to predict amino acid preferences at a residue from its 3D structural environment [26, 27]. It takes as input all atoms within 10Å of a focal residue (with that residue’s atoms masked), featurizing them by atom type, charge, and SASA. HERMES then outputs energies $E_{i}(\sigma)$ for the 20 possible amino acids $\sigma$ at residue $i$, which serve as logits in a softmax to estimate amino acid likelihoods. (B) To predict TCR-pMHC affinity, we use two protocols: HERMES-fixed (left) and HERMES-relaxed (right). In HERMES-fixed we use a fixed the structural template closest to the TCR-pMHC of interest as input to HERMES, and evaluate the peptide energy as the sum of HERMES-computed per-residue energies for the peptide. In HERMES-relaxed mutations are introduced to the template’s peptide to match the peptide of interest, followed by local repacking in PyRosetta [28]. Peptide energy is then evaluated similar to HERMES-fixed, but using the relaxed structure. (C) We also use two protocols for peptide design. Similar to (B) we fix the closest structural template in HERMES-fixed, and use HERMES a likelihood matrix for possible amino acids at each peptide position, which we sample to generate new peptide sequences. In HERMES-relaxed (right), we begin with a random peptide $\sigma$, and locally pack it with PyRosetta, and also generate an amino acid likelihood matrix (similar to HERMES-fixed) for this packed structure. We sample a new sequence $\sigma^{\prime}$ from this likelihood matrix, and pack it with PyRosetta. We evaluate peptide energies $E_{\text{peptide}}(\sigma)$, and $E_{\text{peptide}}(\sigma^{\prime})$ for both peptides, using their respective PyRosetta-relaxed structures, and we use these energies as the criterion to sample improved sequences with MCMC (SI). This procedure is then iterated until convergence. Peptides designed by both approaches are then filtered by TCRdock [17] or Alpha-Fold 3 [29] to assure the stability of the complex.
@@
Here, we present a structure-based approach to predict interaction affinities between TCRs and peptides presented on MHC class I, and to design reactive peptides for specific TCR-MHC-I complexes (TCR-MHC for short). Due to the limited availability of structural data for TCRs, we use HERMES– a physics-guided, 3D-equivariant machine learning model trained on the entire protein universe to predict amino acid preferences based on their local structural environments [27, 26]. We demonstrate that HERMES reliably predicts interaction affinities for diverse TCR-pMHC complexes, and the activity of T-cells in response to peptides presented by MHC class I molecules. Moreover, we show that HERMES can be used to design novel peptides binding specific TCR-MHC complexes, achieving up to 50% experimental validation accuracy across diverse TCR-MHC systems. Our peptide design platform can facilitate the computational design of neoantigen libraries for cancer vaccines. By leveraging our design algorithms, we characterize the specificity of a diverse range of TCRs, providing a quantitative measure for the diversity of the peptide-MHC antigens that a TCR can recognize in humans and mice.
Figure 2: Predicting binding affinity of TCR-pMHC complexes. (A) The structure of the NY-ESO-1 peptide (purple), in complex with HLA-A*02:01 and a purified specific 1G4 TCR is shown (PDB ID: 2BNQ [34]). This structure together with that of a peptide variant with C9V mutation (PDB ID: 2BNR [34]) are used as templates to predict the binding affinity of the TCR-MHC complex to other peptide variants. (B) The binding affinities ($-\log_{10}(K_{d}/\mu\text{M})$) TCR-MHC complexes with 14 distinct peptides, measured using surface plasmon resonance, are shown against the predicted HERMES-relaxed peptide energies, using the closest structural template for each variant; 2BNQ is used for blue points and 2BNR for the orange point. The Spearman $\rho$ between predictions and measurements is reported in the panel. (C) The heatmaps show the experimental affinity values (top) and predicted peptide energies (bottom) for all the peptide variants, with the identity of the mutations color coded based on the closest structural template. (D-F) Similar to (A-C) but for the Tax viral peptide in complex with HLA-A*02:01 and a specific human TCR. Two structural templates are used [35] for evaluating peptide energies with PDB IDs: 1AO7 (shown in D, and used for blue points), and 1QSF (used for orange points).
II Model
@@
Figure 3: Predicting T-cell activity in response to peptide variants. (A-C) The predicted HERMES-fixed peptide energies versus experimentally measured T-cell activity are shown for (A) HLA-A*02:01-restricted CMV peptide variants interacting with TCR1-TCR3 (panels), (B) HLA-A*02:01-restricted melanoma gp100 peptide variants interacting with TCR4-TCR6, and (C) HLA-B*27:05-restricted pancreatic cancer peptide variants interacting with TCR7. In each case, T-cell activity in response to all single amino acid mutations from the wild-type is measured. T-cell activity is based on fitted EC50 values from dose-response curves of 4-1BB+ CD8+ T-cells (SI; data from [36]). Gray points indicate peptides with T-cell responses below detection limit (no EC50 obtained) and are excluded from the reported Spearman correlations. Except for TCR1 (resolved structure, PDB ID: 5D2N [37]), the other TCR-pMHC complexes were modeled with AF3, and the model’s peptide PAE for each wild-type peptide is denoted. These structures are used as templates to estimate peptide energies with HERMES-fixed; see Tables S1, S3 for details on the templates. For all structures, TCR-$\alpha$ (TRA) is shown in blue, TCR-$\beta$ (TRB) in orange, the peptide in magenta, and HLA in gray. (D) The predicted HERMES-fixed peptide energies are shown against experimentally measured T-cell activity for single point mutants of the HLA-A*02:01-restricted p53R175H neoepitope from ref. [5]. T-cell response (IFN-$\gamma$ production) is mediated through a bi-specific antibody binding to the pMHC complex. The co-crystallized structure of the bi-specific antibody with the p53R175H-HLA-A*02:01 complex (PDB ID: 6W51 [5]) is used as the template for our HERMES-fixed predictions. For all systems (A-D), the heatmaps for the predicted and the experimentally measured mutational effects are shown in Fig. S2.
@@
  • HERMES-relaxed: Since amino acid substitutions can locally reshape a protein complex, we introduce a more involved protocol to account for these structural changes. We begin with the closest available structure for the TCR-pMHC complex. We then mutate the original peptide to the desired state and apply a relaxation procedure using PyRosetta to properly pack the substituted amino acid [28] (SI). During relaxation, side-chain and backbone atoms of the peptide are allowed to move, while only the side-chain atoms of the MHC and TCR chains within a 10 Å radius of the peptide are flexible. We then calculate the peptide energy with HERMES (eq.1), using the relaxed structure as input (Fig. 1B); since PyRosetta’s relaxation procedure is stochastic, we average the peptide energy across 100 realizations of these relaxations.
III Results
@@
In our first example, we examined TCR 1G4 binding to NY-ESO-1 peptide variants, which is a cancer-testis antigen commonly expressed in many cancers and is targeted by immunotherapies [45]. We used two structural templates, one with the wild-type peptide SLLMWITQC (NY-ESO-1), and the other with a more immunogenic peptide, in which the Cysteine at position 9 is substituted by Valine (C9V) [46, 34] (Fig. 2A). For variants differing by a single amino acid, our predictions with HERMES-relaxed correlated well (72% Spearman correlation) with experimentally measured affinities from refs. [43, 44](Fig. 2B, C).
@@
Next, we tested the A6 TCR binding to the HTLV-1 Tax peptide LLFGYPVYV. We employed two structural templates: one with the wild-type, the other with the mutant peptide with substitution Y8A [35] (Fig. 2D and Table S1). Again, the HERMES-relaxed peptide energies correlated well (56% Spearman $\rho$) with the affinities, measured in ref. [44] (Fig. 2E, F).
Overall, HERMES predictions aligned closely with this limited experimental affinity measurements and performed as well as or better than BLOSUM-based scoring [47], ProteinMPNN [30], and TCRdock [17] (Table S2). Notably, unlike genome-averaged methods such as BLOSUM, HERMES is sensitive to the local structural environment, assigning different scores to the same mutation depending on context (Fig. S1). HERMES can predict relative peptide binding affinities for TCR-MHC complexes, provided a suitable structural template is available. Prediction quality suffers in the absence of reliable structural templates, as seen also in other methods such as TCRdock [17] (Table S2). Lastly, HERMES-relaxed peptide energies show better agreement with the affinity measurements than HERMES-fixed (Table S2). However, the small size of this dataset prevents us from drawing broader conclusions about the relative utility of either approach for this problem.
@@
Aside from the affinity of TCR-pMHC interactions, other factors play important roles in determining whether and how a T-cell responds. These factors include the phenotypic state of a T-cell (e.g., naive or memory), the reactivity of its TCR to competing self-antigens presented on MHC molecules [48, 49], and the avidity of TCR-pMHC interactions, determined by the clustering of TCRs with co-receptors on the T-cell membrane and the density of pMHC molecules on antigen-presenting cells [50]. T-cell activation can be characterized using a diverse set of functional assays that measure T-cell proliferation, labeled cellular activation markers through flow cytometry, and production of cytokines such as interferon gamma (IFN-$\gamma$). As such, measuring T-cell activity is much easier than measuring the affinity of TCR-pMHC complexes, which requires techniques like SPR. Although our structure-based HERMES model is more attuned to predicting the binding affinity of TCR-pMHC complexes, the ease of data production prompted us to assess the accuracy of our method in predicting T-cell activity.
@@
With the exception of TCR1 (where a TCR-pMHC template with HLA-A*02:01 and one peptide variant was already available), we relied on AlphaFold3 (AF3) [29] to model the remaining TCR-pMHC complexes; see Tables S1, S3 for template information and AF3 sequence inputs for TCRs 2-7. As shown in Figs. 3A, and S2, HERMES-fixed peptide energy predictions (eq. 1) correlate well with experimental EC50, reaching up to 71% Spearman correlation for TCR2 recognizing the CMV peptides. Correlations are lower for TCRs 4-6 targeting variants of the melanoma gp100 self-antigen, with TCR6 achieving 42% Spearman correlation as the best prediction in this class (Fig. 3B). HERMES underperforms for TCR7 responding to the pancreatic cancer neo-peptide GRLKALCQR (Fig. 3C). Still, except for TCR4, which has the largest AF3 Predicted Alignment Error (PAE) of 2.07 at the TCR-pMHC interface (i.e., a lower confidence in its fold), HERMES predictions outperform those of TULIP [15], the state-of-the-art sequence-based language model trained to predict TCR-pMHC interactions (Table S4). The reduced accuracy of HERMES for cancer epitopes compared to the variants of the viral CMV epitope could reflect the impact of antagonism from self-antigens similar to tumor epitopes, hindering T-cell responses [48, 49]—though additional experiments are needed to confirm this.

### v2

@@ Abstract
T-cells play a key role in adaptive immunity by mounting specific responses against diverse pathogens. An effective binding between T-cell receptors (TCRs) and pathogen-derived peptides presented on Major Histocompatibility Complexes (MHCs) mediate an immune response. However, predicting these interactions remains challenging due to limited functional data on T-cell reactivities. Here, we introduce a computational approach to predict TCR interactions with peptides presented on MHC class I alleles, and to design novel immunogenic peptides for specified TCR-MHC complexes. Our method leverages HERMES, a structure-based, physics-guided machine learning model trained on the protein universe to predict amino acid preferences based on local structural environments. Despite no direct training on TCR-pMHC data, the implicit physical reasoning in HERMES enables us to make accurate predictions of both TCR-pMHC binding affinities and T-cell activities across diverse viral epitopes and cancer neoantigens, achieving up to 0.72 correlation with experimental data. Leveraging our TCR recognition model, we develop a computational protocol for de novo design of immunogenic peptides. Through experimental validation in three TCR-MHC systems targeting viral and cancer peptides, we demonstrate that our designs—with up to five substitutions from the native sequence—activate T-cells at success rates of up to 50%. Lastly, we use our generative framework to quantify the diversity of the peptide recognition landscape for various TCR-MHC complexes, offering key insights into T-cell specificity in both humans and mice. Our approach provides a platform for immunogenic peptide and neoantigen design, as well as for evaluating TCR specificity, offering a computational framework to inform design of engineered T-cell therapies and vaccines.
@@ I Introduction
Predicting the TCR-pMHC binding specificity is an important step in characterizing the efficacy of the immune system in countering different pathogens. Such understanding would aid with disease diagnoses and development of diagnostic tests for early detection of autoimmune diseases [1]. A predictive model for TCR-pMHC specificity can be used to design engineered TCRs that recognize cancer antigens [2], enhancing the effectiveness of adoptive cell transfer and Chimeric Antigen Receptor (CAR) T-cell therapies [3, 4], as well as the design of soluble TCRs and bispecific T-cell engager therapeutics [5]. Moreover, it can enable the design of optimized antigens for vaccines to elicit robust T-cell responses against specific pathogens or tumors [6, 7, 8].
@@
One bottleneck in deciphering the TCR-pMHC code is the lack of well-curated datasets. High-throughput immune repertoire sequencing, combined with comparative analyses of hosts facing similar immune challenges, and the development of functional assays (e.g., pMHC-tetramer staining [9] or MIRA [10]), have enabled the identification of diverse TCR pools with potential reactivity to specific antigens [11], collected in repositories like Immune Epitope Database (IEDB) [12] and VDJdb [13]. However, the current databases are unbalanced, with only a handful of dominant peptides (e.g. derived from SARS-CoV-2, CMV, EBV, Influenza, etc.) and MHC alleles present, each linked to hundreds of TCRs.
@@
As a result, statistical inference and machine learning methods trained on such data to predict TCR-pMHC reactivities lack generalizability beyond their training data [14, 15, 16]. On the structural side, only a few hundred TCR-pMHC crystal structures have been experimentally generated to date. This is in part because TCR-pMHC interactions are typically of low affinity and are difficult to stabilize to the level needed for CryoEM or X-ray crystallography [17]. Importantly, this scarcity hinders computational prediction methods like AlphaFold in modeling reliable TCR-pMHC structures [18, 19]. Machine learning models, such as large language models, pre-trained on vast unannotated protein datasets can generate meaningful data representations [20, 21]. Fine-tuning these models on smaller, labeled datasets for specific tasks (e.g., predicting mutation impacts on stability or function) has proven highly effective [22, 23, 24, 25]. Pre-trained sequence-based protein language models have been employed to predict reactivities of antibodies to antigens [24, 25], or TCRs to pMHC complexes [26, 16]. However, lack of generalizability due to sparse and skewed training data has remained a challenge in these models.
@@
Figure 1: HERMES model for predicting and designing TCR-pMHC interactions. (A) HERMES is a structure-based, rotationally equivariant neural network pre-trained on the protein universe to predict amino acid preferences at a residue from its 3D structural environment [27, 28]. It takes as input all atoms within 10Å of a focal residue (with that residue’s atoms masked), featurizing them by atom type, charge, and SASA. HERMES then outputs energies $E_{i}(\sigma)$ for the 20 possible amino acids $\sigma$ at residue $i$, which serve as logits in a softmax to estimate amino acid likelihoods. (B) To predict TCR-pMHC affinity, we use two protocols: HERMES-fixed (left) and HERMES-relaxed (right). In HERMES-fixed we use a fixed structural template closest to the TCR-pMHC of interest as input to HERMES, and evaluate the peptide energy as the sum of HERMES-computed per-residue energies for the peptide. In HERMES-relaxed mutations are introduced to the template’s peptide to match the peptide of interest, followed by local repacking in PyRosetta [29]. Peptide energy is then evaluated similar to HERMES-fixed, but using the relaxed structure. (C) We also use two protocols for peptide design. Similar to (B) we fix the closest structural template in HERMES-fixed, and use HERMES’ likelihood matrix for possible amino acids at each peptide position, which we sample to generate new peptide sequences. In HERMES-relaxed (right), we begin with a random peptide $\sigma$, and locally pack it with PyRosetta, and also generate an amino acid likelihood matrix (similar to HERMES-fixed) for this packed structure. We sample a new sequence $\sigma^{\prime}$ from this likelihood matrix, and pack it with PyRosetta. We evaluate peptide energies $E_{\text{peptide}}(\sigma)$, and $E_{\text{peptide}}(\sigma^{\prime})$ for both peptides, using their respective PyRosetta-relaxed structures, and we use these energies as the criterion to sample improved sequences with MCMC (SI). This procedure is then iterated until convergence. Peptides designed by both approaches are then filtered by TCRdock [18] or Alpha-Fold 3 [30] to assure the stability of the complex.
@@
Here, we present a structure-based approach to predict interaction affinities between TCRs and peptides presented on MHC class I, and to design reactive peptides for specific TCR-MHC-I complexes (TCR-MHC for short). Due to the limited availability of structural data for TCRs, we use HERMES– a physics-guided, 3D-equivariant machine learning model trained on the entire protein universe to predict amino acid preferences based on their local structural environments [28, 27].
We demonstrate that HERMES accurately predicts both binding affinity of TCR-pMHC Class I complexes and the functional responses and T-cell activities induced by a range of peptides presented on MHC molecules. Affinity of TCR-pMHC complexes, measured as the dissociation constant $K_{d}$ by surface plasmon resonance, is labor intensive to obtain. Functional readout used to measure T-cell activities, including measurements of T-cell proliferation, labeled cellular activation markers, and cytokine (e.g., IFN-$\gamma$) release, are easier to obtain than affinity measurements. Yet, functional readouts are shaped not only by the TCR-pMHC affinity, but also by other cellular factors such as the T-cell’s differentiation state (e.g., naive or memory), self-reactivity of the TCR [35, 36], and the avidity effects driven by TCR/co-receptor clustering and pMHC density on antigen-presenting cells [37]. Although HERMES is more suited to predict molecular affinities, we show that its scores nonetheless can well track T-cell activities, making it a reliable tool for predicting T-cell response to antigens.
Lastly, we show that HERMES can be used to design novel peptides binding specific TCR-MHC complexes, achieving up to 50% experimental validation accuracy across diverse TCR-MHC systems. By leveraging our design algorithm, we characterize the specificity of a diverse range of TCRs, providing a quantitative measure for the diversity of the peptide-MHC antigens that a TCR can recognize in humans and mice.
Figure 2: Predicting binding affinity of TCR-pMHC complexes. (A) The structure of the NY-ESO-1 peptide (purple), in complex with HLA-A*02:01 and a purified specific 1G4 TCR is shown (PDB ID: 2BNQ [38]). This structure together with that of a peptide variant with C9V mutation (PDB ID: 2BNR [38]) are used as templates to predict the binding affinity of the TCR-MHC complex to other peptide variants. (B) The binding affinities ($\log_{10}(K_{d}/\mu\text{M})$) of TCR-MHC complexes with 14 distinct peptides, measured using surface plasmon resonance, are shown against the predicted HERMES-relaxed peptide energies, using the closest structural template for each variant; 2BNQ is used for blue points and 2BNR for the orange point. The Spearman $\rho$ between predictions and measurements is reported in the panel. (C) The heatmaps show the experimental affinity values (top) and predicted peptide energies (bottom) for all the peptide variants, with the identity of the mutations color coded based on the closest structural template. (D-F) Similar to (A-C) but for the Tax viral peptide in complex with HLA-A*02:01 and a specific human TCR. Two structural templates are used [39] for evaluating peptide energies with PDB IDs: 1AO7 (shown in D, and used for blue points), and 1QSF (used for orange points).
Model
@@
  • HERMES-relaxed: Since amino acid substitutions can locally reshape a protein complex, we introduce a more involved protocol to account for these structural changes. We begin with the closest available structure for the TCR-pMHC complex. We then mutate the original peptide to the desired state and apply a relaxation procedure using PyRosetta to pack the substituted amino acid [29] (SI). During relaxation, side-chain and backbone atoms of the peptide are allowed to move, while only the side-chain atoms of the MHC and TCR chains within a 10 Å radius of the peptide are flexible. We then calculate the peptide energy with HERMES (eq.1), using the relaxed structure as input (Fig. 1B). Since PyRosetta’s relaxation procedure is stochastic, we average the peptide energy across 100 realizations of these relaxations. For ablation purposes, we also include results obtained by selecting the relaxed conformation with the lowest (most favorable) Rosetta energy, and using its HERMES peptide energy.
Results
@@
TCR-pMHC affinity T-cell activity
Model type Model 1G4 TCR A6 TCR H2-scDb TCR1 TCR2 TCR3 TCR4 TCR5 TCR6 TCR7
Substitution Matrix BLOSUM62 [40] 0.83 0.56* -0.05 -0.07 0.14 0.38 0.33 0.42* 0.55 -0.01
Structure Prediction TCRdock [18] 0.62* 0.88 N/A 0.02 0.13 0.21 0.18* -0.02 0.19 0.06
TCRdock-no template [18] 0.25 -0.14 N/A -0.03 0.44 0.11 -0.22 0.28* 0.14 0.06
Sequence based ML TAPIR [41] 0.29 0.35 N/A 0.36 0.36 -0.00 0.12 0.39* 0.08 -0.01
Structure based ML ESM-IF1 [42] 0.38* 0.17 0.12 0.43* 0.41 0.19 -0.15 0.28* 0.25 0.08
ProteinMPNN [31] -0.01 0.49 0.55 0.46* 0.52 0.49* -0.28 0.45 0.41* 0.31
Structure based ML HERMES-fixed 0.29 0.63* 0.46* 0.56 0.71 0.59 -0.06 0.37* 0.42* -0.01
HERMES-relaxed 0.72* 0.56* 0.29 0.31 0.57 0.55* -0.20 0.37* 0.42* 0.14*
Table 1: Benchmarking of models for predicting TCR-pMHC binding affinities and T-cell activities. The table lists the Spearman $\rho$ of models’ predictions against experimentally measured binding affinities (Fig. 2) and T-cell activities (Fig. 3) in response to peptide variants. The HERMES-fixed and HERMES-relaxed performance are from models with no noise (0.00) in training data, and the ProteinMPNN performance is from the model with a small noise amplitude of 0.02 Å; see SI for details. We show in bold the best model for each dataset, and any model whose performance is not significantly different from the best model (p-value>0.05>0.05 associated with the difference in the Fisher’s z-transformed Spearman correlations) is indicated with an asterisk (*). See Tables S2 and S3 in the SI for $p$-values and additional models.
@@
In our first example, we examined TCR 1G4 binding to NY-ESO-1 peptide variants, which is a cancer-testis antigen commonly expressed in many cancers and is targeted by immunotherapies [50]. We used two structural templates, one with the wild-type peptide SLLMWITQC (NY-ESO-1), and the other with a more immunogenic peptide, in which the Cysteine at position 9 is substituted by Valine (C9V) [51, 38] (Fig. 2A). For variants differing by a single amino acid, our predictions with HERMES-relaxed correlated well (Spearman $\rho=0.72$) with experimentally measured affinities from refs. [48, 49]; see Fig. 2B, C.
@@
Next, we tested the A6 TCR binding to the HTLV-1 Tax peptide LLFGYPVYV. We employed two structural templates: one with the wild-type, the other with the mutant peptide with substitution Y8A [39] (Fig. 2D and Table S1). Again, the HERMES-relaxed peptide energies correlated well (Spearman $\rho=0.56$) with the affinities, measured in ref. [49] (Fig. 2E, F).
Overall, HERMES predictions aligned closely with these limited experimental affinity measurements, with HERMES-relaxed showing a better agreement with affinity measurements than HERMES-fixed. We compared our approach for scoring peptides to alternative methods, including the baseline substitution matrices (BLOSUM62 [40] and the TCR-specific substitution matrices from ref. [52]), the score from the TCR-pMHC structure prediction algorithm TCRdock [18], the sequence-based machine learning algorithms for TCRs (TAPIR [41], NetTCR [53], and TULIP [16]), and the structure-based inverse folding models (ESM-IF1 [42] and ProteinMPNN [31]); see Table 1 for key benchmarking results and Table S2 and Fig. S1 for a more extensive comparison of different methods, including the different HERMES models, and different scoring schemes by ProteinMPNN.
For the 1G4 TCR, BLOSUM62 [40] achieves the highest correlation with experimental affinity data, whereas template-guided TCRdock [18] leads for the A6 system; in contrast, TCRdock without using a homologous structure as a template underperforms. In both systems, HERMES achieves performance that is statistically indistinguishable from the top performers (Fisher’s z-test p-value >0.05>0.05). Overall, BLOSUM62, template-guided TCRdock, and HERMES rank strongly as predictors of TCR-pMHC affinities across the tested datasets. In addition, for the 1G4 system, ESM-IF1 [42] and the TCR-specific substitution matrices from ref. [52] also perform competitively. However, due to the limited number of affinity measurements, we lack statistical power to resolve finer differences among the models.
It should be noted that although BLOSUM62 is one of the best contenders for affinity prediction, its sequence‑averaged construction ignores local structural context and therefore assigns identical substitution scores regardless of a residue’s environment (Fig. S2); similar limitations apply to the TCR-specific substitution matrices from ref. [52]. HERMES, by contrast, is structure‑aware and can attribute distinct scores to the same mutation when it occurs in different conformational settings–an ability that is critical for nuanced mutational forecasts (Fig. S2). Determining the biophysical circumstances under which such context‑agnostic scoring schemes suffice, and when structure‑aware approaches become necessary, will require systematic evaluation on broader, more diverse datasets.
Lastly, we probed HERMES’s robustness to potential inaccuracies in structures generated by the widely used AlphaFold3 (AF3) algorithm [30], replacing the crystal structure inputs with AF3 models; see SI for details. Both the template-guided and the de novo AF3 structures remained within $1.1$ Å  RMSD of the respective crystal structures (Table S4). For both HERMES and ProteinMPNN, replacing experimental structures with AF3 predictions altered predictions for the 1G4 and A6 TCRs binding to pMHC variants, but the changes were not statistically significant (Fisher’s z-test p-value >0.05>0.05); see Fig. S4. It is nonetheless expected that the predictions should suffer in the absence of reliable structural templates, as seen also in other methods such as TCRdock [18] (Tables 1, S2).
Figure 3: Predicting T-cell activity in response to peptide variants. (A-C) The predicted HERMES-fixed peptide energies versus experimentally measured T-cell activity are shown for (A) HLA-A*02:01-restricted CMV peptide variants interacting with TCR1-TCR3 (panels), (B) HLA-A*02:01-restricted melanoma gp100 peptide variants interacting with TCR4-TCR6, and (C) HLA-B*27:05-restricted pancreatic cancer peptide variants interacting with TCR7. In each case, T-cell activity in response to all single amino acid mutations from the wild-type is measured. T-cell activity is based on fitted EC50 values from dose-response curves of 4-1BB+ CD8+ T-cells (SI; data from [52]). Gray points indicate peptides with T-cell responses below detection limit (no EC50 obtained) and are excluded from the reported Spearman correlations. Except for TCR1 (resolved structure, PDB ID: 5D2N [54]), the other TCR-pMHC complexes were modeled with AF3, and the model’s confidence for each wild-type peptide is denoted. These structures are used as templates to estimate peptide energies with HERMES-fixed; see Table S1 and Dataset S1 for details on the templates. In all panels, TCR-$\alpha$ (TRA) is shown in blue, TCR-$\beta$ (TRB) in orange, the peptide in magenta, and HLA in gray. (D) The predicted HERMES-fixed peptide energies are shown against experimentally measured T-cell activity for single point mutants of the HLA-A*02:01-restricted p53R175H neoepitope from ref. [5]. T-cell response (IFN-$\gamma$ production) is mediated through a bispecific antibody binding to the pMHC complex. The co-crystallized structure of the bispecific antibody with the p53R175H-HLA-A*02:01 complex (PDB ID: 6W51 [5]) is used as the template for HERMES-fixed predictions. For all systems (A-D), the heatmaps for the predicted and the experimentally measured mutational effects are shown in Fig. S3.
@@
T-cell activation is shaped by factors beyond TCR-pMHC affinities, including the T-cell’s differentiation state, its cross-reactivity to self-antigens [35, 36], and avidity effects from TCR/co-receptor cross-linking [37]. Nonetheless, because readouts of T-cell activiy (e.g., T-cell proliferation, and production of activation markers) are easier to obtain than SPR-based affinity measurements, we next asked whether HERMES scores, which are more attuned for predicting TCR-pMHC binding affinity, can serve as predictors of T-cell activities.
@@
With the exception of TCR1 (where a TCR-pMHC template with HLA-A*02:01 and one peptide variant was already available), we relied on AlphaFold3 (AF3) [30] to model the remaining TCR-pMHC complexes; see Table S1 and Dataset S1 for template information and AF3 sequence inputs for TCRs 2-7. As shown in Figs. 3A, and S3, HERMES-fixed peptide energy predictions (eq. 1) correlate well with experimental EC50 measurements for the CMV peptide variants, reaching up to Spearman correlation $\rho=0.71$ for TCR2, and outperforming the alternative models. For TCR1 and TCR3, the performance of ProteinMPNN [31] is worse but statistically indistinguishable from HERMES (Tables 1, S3, and Fig. S1). Moreover, replacing the crystal structure for TCR1 with an AF3 model yields comparable performance, underscoring HERMES’s robustness to structural inputs so long as the structural models are of high quality (Figs. S5, S4). Correlations are lower for TCRs 4-6 targeting variants of the melanoma gp100 self-antigen, with HERMES achieving Spearman correlation $\rho=0.42$ for TCR6 as its best prediction in this class, with an accuracy comparable to ProteinMPNN [31] (Fig. 3B and Tables 1, S3). HERMES underperforms for TCR7 responding to the pancreatic cancer neo-peptide GRLKALCQR (Fig. 3C), for which only the ProteinMPNN’s predictions show significant albeit weak correlation with data (Table 1). The reduced accuracy of HERMES for cancer epitopes compared to the variants of the viral CMV epitope could reflect the impact of antagonism from self-antigens similar to tumor epitopes, hindering T-cell responses [35, 36]—though additional experiments are needed to confirm this.
Sequence-based models (TAPIR [41], NetTCR [53], and TULIP [16]) are outperformed by structure-based models tested here in all cases except for TCR4, which has the lowest AF3 confidence at the TCR-pMHC interface (i.e., a lower confidence in its fold); see Table S4 and Fig. S4 for the AF3 confidence metrics for all systems. Here, even the basic substitution models (BLOSUM62 [40] and the TCR-specific matrices from ref. [52]) outperform the structure-based predictors (Tables 1, S3). This suggests that structural approaches hinge on a reliable structure as input, though more data is needed to draw definitive conclusions (Fig. S4).